OpenStudy (anonymous):

Tonya is twice Kevin's age. In 3 years, Tonya will be 17. How old is Kevin?

OpenStudy (anonymous):

Kevin is like 7 i think....

OpenStudy (anonymous):

Subtract 3 years, 17-3=14 14/2 = 7

OpenStudy (anonymous):

7 Sorry, i misread your question

OpenStudy (anonymous):

i think its like this

OpenStudy (anonymous):

2(k+3)=17 to get 2k+6=17

OpenStudy (anonymous):

kevin is 5.5 years old, tonya is 11. in 3 years kevin will be 8.5 yo and tanya will be 2(8.5)=17

OpenStudy (anonymous):

Kevin is 7. In 3 years, Tonya will be 17. 17-3 is 14. She is twice Kevin's age. 14/2 is 7.